On the first 5 quizzes of his physics class, Daniel got an average score of 85.
What does he need on the next quiz to have an overall average of 87?

Answer by Cromlix(4381)

5 quizzes average score = 85
Total scores = 85 * 5 = 425
Total scores = 87 * 6 = 522
So Daniel requires a score of(522 - 455)
= 67 to attain an average of 87 with
the 6 quizzes.
Hope this helps.
:-)